OpenStudy (anonymous):

what is the particular solution for this differential equation y''''+y''=x^4

OpenStudy (anonymous):

try let y=Ax^4+Bx^3+Cx^2+Dx+E

OpenStudy (anonymous):

i did that sir it does not work

OpenStudy (anonymous):

all the coefficients can not be solved for by using that polynomial

OpenStudy (anonymous):

i would prefer u http:/www.wolframalpha.com

OpenStudy (nilankshi):

hmmmmm

OpenStudy (anonymous):

yp=ax^6+bx^5+cx^4+dx^3+ex^2+fx+g

OpenStudy (anonymous):

@ ictrees you must be at school already! get to class

OpenStudy (anonymous):

Still say set y'' = z so y'''' = z'', make life simpler...

OpenStudy (anonymous):

u sayin use laplace?

OpenStudy (anonymous):

No, u just looking for generic solution, not ivp, right?

OpenStudy (anonymous):

yea...just generic i need to know what method i need to use...

OpenStudy (nilankshi):

ohh

OpenStudy (anonymous):

why ohhh?

OpenStudy (nilankshi):

nothingohhh

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So to get the complementary u solve the auxiliary (for the reduced equation) I think...

OpenStudy (anonymous):

i used yp = Ax^4+Bx^3+Cx^2+Dx+E, but the Dx part is repeated from complementary solution so you have to multiply it by X^2 to get rid of that repeating solution..i even plugged my solution in and checked it worked out perfect

OpenStudy (anonymous):

for those that are curious the full solution is y=c1+c2x+c3cosx+c4sinx+(1/30)(x^6)-(x^4)+(12)(x^2)

OpenStudy (anonymous):

I see complementary (for reduced) is just c1 sin + c2 cos (lambda = +-i) and Wolfram then gives x^4-12x^2 +24 then you can integrate twice to get back to your equation.

OpenStudy (anonymous):

what is wolfram?

OpenStudy (anonymous):

btw estudier this is not a second order equation its a fourth order equation and other term is second order

OpenStudy (anonymous):

Sure but u can set z = y'' to convert to second order then just integrate twice at the end.

OpenStudy (anonymous):

U can see it comes out the same, no?

OpenStudy (anonymous):

no sir...

OpenStudy (anonymous):

Why not...?

OpenStudy (anonymous):

the zero solutions are missed out if u transform the equation

OpenStudy (anonymous):

The solution u get is the same as the solution u gave....

OpenStudy (nilankshi):

i cannt uderstand the qus also :(

OpenStudy (anonymous):

is it1/30 x^4-x^2-12